what is the largest rectangular area that can be enclosed with 16 meters of fence?
Answer by nerdybill(7384)

what is the largest rectangular area that can be enclosed with 16 meters of fence?
Let L = length
and W = width
.
2(L + W) = 16 (equation 1)
area = WL (equation 2)
.
Solve equation 1 for L:
2(L + W) = 16
(L + W) = 8
L = 8-W
.
Substitute above into equation 2:
area = W(8-W)
area = 8W-W^2
area = -W^2+8W
this is a quadratic (parabola) that opens downwards (from the negative coefficient associated with the W^2 term). This means the vertex is the Max.
max is at
W = -b/(2a)
W = -8/(2*(-1))
W = -8/(-2)
W = 4 meters (width)
.
L = 8-W = 8-4 = 4 meters (length)
.
Maximum area is
WL = 4*4 = 16 square meters